OpenStudy (anonymous):

Solve for x and y. x + 4 = 2y x – 7 = 3

OpenStudy (acxbox22):

might as well use yahoo for answers

OpenStudy (anonymous):

is there no y in the second one?

OpenStudy (anonymous):

\[x + 4 = 2y\\ x – 7 = 3\] or \[x + 4 = 2y \\x – 7y = 3\]

OpenStudy (anonymous):

x-7=3 add 7 to both sides: x=10 x + 4 = 2y substitute 10 in where x is: 10+4=2y so,14=2y then, divide both sides by 2 7=y

OpenStudy (anonymous):

so x=10 and y=7